Fermi National Laboratory


Accelerator Update - July 7 - July 9

The integrated luminosity for last week was 2719.1 nb-1.

Monday July 7
The day shift began with the Antiprotons Source (Pbar) stacking, and with beam to MiniBooNE.

Operations put the Tevatron (TeV) into shot setup at 8:29 AM.

At 9:50 AM, a Pbar Debuncher power supply (D:VT925) stopped putting out current, which delayed the shot setup. Operators contacted an expert who had the supply operating normally by 10:13 AM.

Operations established store 2768 at 11:31 AM with an initial luminosity of 31.06E30. Pbar resumed stacking.

Operations reported seeing a possible correlation, at 1:15 PM, between the rise in outside temperatures and Booster MH2 power supply problems. Its regulation suffered whenever line voltage drooped.

Cryo technicians at CHL reported that the wet engine for TeV sector C3 was having problems. The cryo coordinator did not believe it will last through the night. Experts were called in to investigate.

The Main Injector (MI) power supplies at MI-10 tripped off at 9:44 PM. Operators called an expert who instructed the operators to check certain systems. Nothing of any significance was found. Stacking resumed at 10:05 PM.

A cryo system expert investigated the TeV sector C3 wet engine, at 11:19 PM, and confirmed it would not last much longer. If this engine was turned off or tripped off the store would be lost. They were devising a plan to try using a JT valve to keep the flow of helium to the cryo magnets when a power glitch, at 11:20 PM, caused the store to abort. Too many things tripped off to bother naming. Everything was, eventually, turned back on.

The Collider Detector Facility (CDF) and D-zero (D0) experiments accessed their collision halls at 11:36 PM.

Pbar resumed stacking at 11:42 PM.

Work on the TeV sector C3 wet engine began at 11:50 PM.

Tuesday July 8
The midnight shift began with Pbar stacking, with MiniBooNE taking beam, and with the TeV recovering from the power glitch.

D0 came out of access at 12:31 AM, and CDF came out an hour later.

The cryo techs reported, at 2 AM, that the C3 wet engine needed its flywheel changed. They completed their work at 3:14 AM.

Operations put the TeV through a dry squeeze and then began shot setup at 3:49 AM.

Operations established store 2770 at 6 AM. Its initial luminosity was 32.37E30. Pbar resumed stacking.

At 8:58 AM, Operations turned off the Linac beam so an expert could work on the Booster Debuncher. Beam returned ten minutes later.

A Booster RF station (BRF9) tripped off seven times between 4:22 PM and 5:43 PM. Operators talked with an expert who advised lowering the modulator's voltage; this stopped the sparking.

MiniBooNE asked to have their beam stopped, at 6:34 PM, so AC techs could repair their detector's air conditioning. They were back on line at 8 PM.

A Recycler expert began a study period at 8:38 PM.

A Booster device (B:CEPSE2) tripped off at 8:42 PM. Operators investigated, found and changed a bad MADC chassis.

Pbar resumed stacking at 9:28 PM.

A Booster kicker (MKS03) tripped off at 9:31 PM. Operations called in two experts who had the kicker working by 11 PM.

At 10:49 PM, a Booster device (B:CEPSE2) tripped again, this time due to power supply fault indication.

Booster suffered from a flurry of RF problems at 11 PM. All the RF stations were back on by 11:19 PM except for station 13. An expert, called in earlier, arrived and began investigations and possible repairs.

Pbar resumed stacking at 11:40 PM.

Wednesday July 9
The midnight shift began with Operations monitoring store 2770, with Pbar stacking, with Switchyard studies, and with MiniBooNE taking beam.

A TeV sector A2 wet engine tripped off at 12:21 AM, due to a gas bottle going empty that operated the engine's brake. A cryo tech changed out the bottle.

The expert investigating the BRF13 problems stated, at 12:30 AM, that the station's modulator needed changing. The station was bypassed and the repair was put off until a later date.

Operations terminated store 2770 at 2:05 AM.

The Switchyard expert finished his study at 2:14 AM.

Operations put the TeV through a dry squeeze at 2:53 AM and then into shot setup at 3:23 AM.

Operations established store 2772 at 7 AM with an initial luminosity of 33.38E30. Pbar resumed stacking.

Today's Plan
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/9/2003   email Fermilab